AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 296     Entry time: Sun Feb 27 05:45:53 2022
Author: David, Pierre-Michell 
Subject: Current status 

6:05: 95K-->120K

6:12: 120K --> 95K

Ring current 8e7 particles/s

Merger, Tape server (WS3): 1414 kB/s writing ok

Leakage current (WS4): 3.25 uA

DAQ (WS2): temperature checked, clock status: ok, adc calibration: ok, white rabbit check: failed

Online monitoring (WS 5): done

Attachment 1: 2022-02-27_06-50-09exrat.png  156 kB  | Hide | Hide all
2022-02-27_06-50-09exrat.png
Attachment 2: 2022-02-27_06-48-23eytot.png  167 kB  | Hide | Hide all
2022-02-27_06-48-23eytot.png
Attachment 3: 2022-02-27_06-48-05extot.png  167 kB  | Hide | Hide all
2022-02-27_06-48-05extot.png
Attachment 4: 2022-02-27_06-47-47xytot.png  174 kB  | Hide | Hide all
2022-02-27_06-47-47xytot.png
ELOG V3.1.4-unknown